### v3

# Accessible, At-Home Detection of Parkinson’s Disease via Multi-task Video Analysis

Abstract
Limited accessibility to neurological care leads to underdiagnosed Parkinson’s Disease (PD), preventing early intervention. Existing AI-based PD detection methods primarily focus on unimodal analysis of motor or speech tasks, overlooking the multifaceted nature of the disease. To address this, we introduce a large-scale, multi-task video dataset consisting of $1102$ sessions (each containing videos of finger tapping, facial expression, and speech tasks captured via webcam) from $845$ participants ($272$ with PD). We propose a novel Uncertainty-calibrated Fusion Network (UFNet) that leverages this multimodal data to enhance diagnostic accuracy. UFNet employs independent task-specific networks, trained with Monte Carlo Dropout for uncertainty quantification, followed by self-attended fusion of features, with attention weights dynamically adjusted based on task-specific uncertainties. To ensure patient-centered evaluation, the participants were randomly split into three sets: $60\%$ for training, $20\%$ for model selection, and $20\%$ for final performance evaluation. UFNet significantly outperformed single-task models in terms of accuracy, area under the ROC curve (AUROC), and sensitivity while maintaining non-inferior specificity. Withholding uncertain predictions further boosted the performance, achieving $88.0\pm 0.3\%$ accuracy, $93.0\pm 0.2\%$ AUROC, $79.3\pm 0.9\%$ sensitivity, and $92.6\pm 0.3\%$ specificity, at the expense of not being able to predict for $2.3\pm 0.3\%$ data ($\pm$ denotes 95% confidence interval). Further analysis suggests that the trained model does not exhibit any detectable bias across sex and ethnic subgroups and is most effective for individuals aged between 50 and 80. Requiring only a webcam and microphone, our approach facilitates accessible home-based PD screening, especially in regions with limited healthcare resources.
Introduction
Due to limited access to neurological care, many individuals, particularly in underserved regions (Kissani et al. 2022), live with Parkinson’s disease—the fastest growing neurological disorder (Dorsey et al. 2018)—without even knowing it. In many instances, when diagnosed late with this incurable disease, the condition could have already progressed significantly, limiting the usefulness of available medications. Imagine a future where individuals can remotely assess their risk for PD by simply visiting a website, activating their webcam and microphone, and completing a series of standardized tasks. This accessible approach could empower people to seek early intervention and treatment, potentially improving their quality of life.
Detecting PD is particularly challenging due to the variability of individual symptoms. For example, while vocal impairment is a common PD symptom (Ho et al. 1999), patients may also exhibit PD through other modalities, such as facial expression (e.g., hypomimia (Gunnery et al. 2016)) or motor function (e.g., bradykinesia (Bologna et al. 2023)). Consequently, models relying solely on a single task may yield suboptimal performance. To address this, we introduce a large video dataset featuring webcam recordings of individuals performing three tasks: (i) finger-tapping (motor function), (ii) smiling (facial expression), and (iii) uttering a pangram111pangram: sentence containing all the letters of the alphabet. (speech). The dataset is collected from 845 unique participants (272 with PD) from diverse demographics who recorded all of these tasks successfully (about 20% of them recorded multiple times), resulting in 1102 videos for each task and a total of 3306 videos. To the best of our knowledge, this is the first multi-task video dataset for PD screening.
We propose a novel two-stage classification model to distinguish between individuals with and without Parkinson’s disease (PD) using our dataset. First, each task is independently modeled with neural networks trained using Monte Carlo dropout (MC-dropout) to generate predictions and associated uncertainties. Next, we introduce the Uncertainty-calibrated Fusion Network (UFNet), which aggregates features from multiple tasks through an attention mechanism (Vaswani et al. 2017), while calibrating the attention scores based on task-specific uncertainties. UFNet is designed to produce PD/Non-PD predictions and uses MC-dropout during both training and inference to quantify prediction confidence. By filtering out low-confidence predictions, it increases patient safety. Our evaluations on a subject-separated test set demonstrate that UFNet significantly outperforms both single-task models and multi-modal fusion baselines. The model is computationally efficient, employing only a few shallow neural networks and a single self-attention module, making it suitable for deployment on smartphones or personal computers.
Here is a summary of our key contributions:
  • We introduce the first-ever large-scale, multi-task video dataset for Parkinson’s disease screening collected from 845 individuals with diverse demography. Although we cannot share the raw videos to protect identifiable patient information, we will release the dataset in the form of extracted features (supplied as supplementary materials).
  • We propose a novel, effective multimodal fusion model named UFNet, achieving $88.0\pm 0.3\%$ accuracy and $93.0\pm 0.2\%$ AUROC on a subject-separated test dataset. The proposed model significantly outperformed single-task models and multimodal fusion baselines.
Related Works
Traditionally, PD is diagnosed by a clinician based on medical history and a clinical examination, which typically includes the completion of standardized tasks and rating each task according to the MDS-UPDRS guidance (Goetz et al. 2008). Recently, a cerebrospinal fluid (CSF) based $\alpha$-synuclein seed amplification assay has been developed (Siderowf et al. 2023), offering a potential diagnostic biomarker. However, access to clinical care can be limited, and diagnostic methods relying on the collection of CSF are invasive, costly, and burdensome to the patients.
Recent research has leveraged machine learning and sensors to facilitate remote assessment of PD. For example, nocturnal breathing signals, obtained from a breathing belt or reflected radio waves, can detect PD with high accuracy when analyzed by a machine learning model (Yang et al. 2022). Additionally, several body-worn sensors have been successfully used to monitor clinical features such as dyskinesia and gait disturbances associated with PD (Moreau et al. 2023). However, wearable sensors face challenges related to cost, comfort, and ease of use, which limit the scalability of these techniques for global adoption.
Using advanced computer vision and machine learning tools, precise clinical features can be extracted from recorded videos and used to screen for PD (Jin et al. 2020) or track symptom progression (Islam et al. 2023). However, existing methods for video analysis suffer from two major limitations – small cohort size (Jin et al. 2020), and reliance on single modality (Tsanas et al. 2009; Ali et al. 2020, 2021; Rahman et al. 2021; Islam et al. 2023; Adnan et al. 2023). Symptoms of PD are multi-faceted and may affect individuals differently. For instance, one individual may face speech difficulty but retain relatively normal motor functionality, while another individual may have prominent hypomimia (i.e., reduced facial expression) or bradykinesia (i.e., slowness incoordination of movements). Therefore, PD detection models may need to consider all these different modalities for improved efficacy.
This work addresses a critical gap in the field, which has persisted due to the lack of multimodal, diverse, and naturalistic data. Traditionally, clinical enterprises often lack the insight or foresight to determine what data to collect and in what quantities, while technical teams are constrained by limited access to patients and resources, preventing them from dedicating years to gathering such comprehensive data. These challenges have largely prevented the development of robust multimodal PD detection models. By establishing a new collaboration, we have been able to overcome these barriers and set new benchmarks for a critical disease. With the features we have made available, we hope the AI community will engage with this dataset and continue to advance the model further.
Dataset
Standardized Tasks
We selected three standardized tasks that can be easily completed using a computer webcam and microphone, with or without external supervision:
  • (i)
    Finger-tapping: participants tap their thumb with the index finger ten times as fast as possible, first with the right hand, then with the left. Finger-tapping task is completed in accordance with the MDS-UPDRS scale to measure bradykinesia in the upper limb, a key sign of PD (Hughes et al. 1992).
  • (ii
    Smile: participants mimic a smile expression three times, alternating with a neutral face. While the expression may be unnatural, studies suggest it still captures signs of hypomimia (Bandini et al. 2017; Adnan et al. 2023).
  • (iii)
    Speech: participants utter a script, “The quick brown fox jumps over a lazy dog. The dog wakes up, and follows the fox into the forest. But again, the quick brown fox jumps over the lazy dog.” The first sentence is a pangram, containing all the letters of the English alphabet, and the later sentences are added to obtain a longer speech segment. Prior research identified this task as a promising way of screening PD (Rahman et al. 2021).
Participant Recruitment
We recruited a diverse cohort of participants (with and without PD) through multiple channels including a brain health study registry, social media, a PD wellness center, and clinician referrals. Among the approximately 1,400 individuals who recorded at least one task, 845 completed all three standardized tasks (comprising this dataset). Participants were primarily from the US, with diverse demographic characteristics: 52.7% female, a mean age of 61.9 years (22.5% over 70), and representation from various ethnic backgrounds. Among the 272 (32.2%) participants who had PD, 233 had their diagnosis confirmed through clinical evaluation, while 39 self-reported their condition. A detailed demographic summary of the participants is available in the supplementary materials (Technical Appendix A). This study was approved by relevant institutional review boards.
Task/Split #Sessions, PD % #Participants, PD %
Finger-tapping 1374, 41.3% 1167, 36.6%
Training 945, 43.9% 819, 38.8%
Validation 221, 37.6% 172, 32.6%
Test 208, 33.2% 176, 30.1%
Smile 1684, 32.8% 1357, 28.5%
Training 1021, 33.2% 824, 28.4%
Validation 342, 33.9% 266, 28.6%
Test 321, 30.5% 267, 28.8%
Speech 1655, 33.9% 1265, 28.9%
Training 1007, 35.3% 769, 29.0%
Validation 338, 33.7% 252, 29.0%
Test 310, 29.7% 244, 28.7%
All tasks $\mathbf{1102,41.8\%}$ 845, 32.2%
Training 690, 45.1% 516, 32.6%
Validation 215, 38.1% 167, 32.9%
Test 197, 34.9% 162, 30.2%
Table 1: The single-task datasets contain videos of three different tasks: finger-tapping (motor function), smiling (facial expression), and speech (pangram utterance) from 1167, 1357, and 1265 participants, respectively. The multi-task dataset comprises 845 unique participants with videos of all three tasks. Validation and test set participants are the same across single-task and multi-task experiments (participants with missing videos are excluded).
Dataset Splits
While most participants completed all three tasks, some did not complete one or more tasks, and some task videos were discarded due to feature extraction failures, often caused by inaccurate task completion. Instead of discarding videos of participants with missing tasks, we trained task-specific models with all available videos for each task. The multi-task model is then trained on the participants who completed all three tasks. We split the datasets based on the participants to ensure patient-centric evaluation. First, we listed all the participants ($n=1402$) enrolled in the study and then randomly assigned 60%, 20%, and 20% of the participants into the training, validation, and test sets, respectively. The random assignment was stratified to ensure a similar ratio of individuals with and without PD across these three folds. Both task-specific and multi-task models are validated and tested on the same participant cohort, with no model seeing any data from these participants during training. Table 1 provides a summary of the dataset splits.
Our Approach
Figure 1: An illustrative overview of the proposed model. Task-specific models are shallow neural networks trained with task-specific features and MC-dropout. Standard deviations in multiple rounds of inference obtained from the task-specific models are used to calibrate the attention scores when fusing task-specific features. Finally, another shallow neural network is trained to differentiate individuals with and without PD, and withhold prediction when the model is uncertain. Authors have obtained consent to publish images of the human subject.
Feature Extraction
We rely on prior literature to extract clinically meaningful features for each task. Although state-of-the-art deep learning models can learn to represent a video without explicit feature extraction (Liu et al. 2022b; Tong et al. 2022), training these models would require a notably larger dataset. By converting task-videos into feature sets, we can use simpler models with fewer trainable parameters.
Finger-tapping features.
Islam et al. (2023) extracted 65 features to analyze the finger-tapping task for assessing PD severity. Using MediaPipe (Grishchenko et al. 2022) hand to detect hand movements and key-points, they measured clinically relevant features such as finger-tapping speed, amplitude, and interruptions. We apply this technique to both hands, extracting 130 features in total.
Smile features.
We leveraged the 42 facial features extracted by Adnan et al. (2023) from smile mimicry videos. These features, captured using OpenFace (Baltrušaitis, Robinson, and Morency 2016) and MediaPipe, encompass key PD markers outlined by the MDS-UPDRS, such as eye blinking, lip separation, mouth opening, and intensity of facial muscle movements.
Speech features.
We extracted 1024-dimensional embeddings from a pre-trained WavLM (Chen et al. 2022) language model to encode the pangram utterance task. WavLM, trained on massive amounts of speech data, excels at capturing acoustic characteristics of speech, making it useful for various tasks like speech recognition, speaker identification, and emotion recognition. Recent research has shown WavLM embeddings to be effective for PD screening (Adnan et al. 2024).
Task-specific Models
Each task utilizes a separate machine learning model to distinguish between individuals with and without PD. These models have three main components:
  • Feature selection and scaling: Pairwise correlation among the features is calculated based on the training data. If two features have a Pearson’s correlation coefficient above a specified threshold, one feature is dropped. Feature values are scaled using StandardScaler or MinMaxScaler. Whether to apply feature selection or scaling, the correlation threshold, and the scaling method are hyper-parameters tuned on the validation set.
  • Shallow neural networks: The shallow neural network consists of 1-2 linear layers, with the last layer having a single output neuron. We use sigmoid activation with the output layer and ReLU with the other layers to add non-linearity.
  • Monte Carlo dropout (MC-dropout): To improve model robustness and estimate prediction uncertainty, we employ MC-dropout (Atighehchian et al. 2022). This technique allows using dropout for both training and inference. Therefore, we can obtain predictions from multiple rounds during testing. After $n$ rounds of inference, we obtain the mean ($\mu$) and standard deviation ($\sigma$) across $n$ different predictions for the same data. When the model is uncertain, we expect a relatively higher standard deviation across these predictions. Therefore, $\sigma$ is used to estimate model uncertainty.
The task-specific models are trained using binary cross-entropy loss, along with an optimizer (see Technical Appendix C regarding hyper-parameter tuning).
Uncertainty-calibrated Fusion Network, UFNet
The fully trained task-specific models remain frozen during the training of UFNet. For each task $i$, the extracted features ($X_{i}\in\mathbb{R}^{d_{X_{i}}}$), predicted mean probability ($\mu_{i}$), and uncertainty in the prediction ($\sigma_{i}$) are input to UFNet. The model then combines information from all the tasks through a series of steps to generate a final, more robust detection of PD (see Figure 1 for an overview).
Projection.
Since the size of the features may vary from task to task, they are first projected to the same dimension ($d$) using a projection layer. Each projection layer consists of a linear layer ($\mathbb{R}^{d_{X_{i}}}\rightarrow\mathbb{R}^{d}$) with MC-dropout, followed by non-linear activation (ReLU) and layer normalization.
Calibrated self-attention.
We employ a self-attention mechanism to integrate the projected features ($X^{p}_{i}\in\mathbb{R}^{d}$) extracted from three distinct tasks. Unlike standard self-attention applications, our input sequence comprises task-specific feature vectors (sequence length = 3) rather than tokens of the same modality. For each task, projected features are transformed into query, key, and value representations using learned linear projections. Attention weights are computed based on the similarity between queries and keys.
To prioritize informative tasks, we adjust the attention scores to down-weight contributions from tasks with higher prediction uncertainty. Specifically, in addition to using traditional query-key similarity, we update the attention matrix $A$ with task-specific uncertainty scores $\Sigma=[\sigma_{1},\sigma_{2},\sigma_{3}]$ as,
$A=softmax(\frac{Q.K^{T}}{\sqrt{d}}-\eta\Sigma)$
Here, $Q,K$ are the query and key matrices respectively, and $\eta$ is a hyper-parameter. After computing the attention scores, $X^{p}_{i}$ is converted into a contextualized representation $Z_{i}=\sum_{j=1}^{N}a^{\prime}_{j}v_{j}$, where $v_{j}$ is the corresponding value vector, and $N=3$ is the number of tasks. Please refer to Technical Appendix E for Pytorch implementation of the self-attention module.
Shallow neural network.
The contextualized representations ($Z_{1},Z_{2},Z_{3}$) obtained after self-attention are concatenated along with the task-specific (mean) predicted probabilities ($\mu_{1},\mu_{2},\mu_{3}$). The merged vector is then input to a shallow neural network similar to the one used for task-specific training. The network is trained with 30 rounds of MC dropout, and the average output is used as the final prediction (PD if the average output is more than 0.50, non-PD otherwise).
Withholding predictions.
Using predictions obtained from multiple rounds of MC-dropout, we can model the confidence of the predictions. Here, we compute the $95\%$ confidence interval of the predicted scores, and if the interval contains the decision threshold (i.e., 0.50), the prediction is considered to be of low confidence. For patient safety, we withhold such predictions as they are more likely to be inaccurate.
Finally, the model is trained with binary cross-entropy loss and SGD optimizer with momentum $0.6898$. After hyper-parameter tuning on the validation set, the query dimension is set to 64 with a $0.0207$ learning rate, $0.4960$ dropout probability, and $\eta=81.8179$ (see Technical Appendix D for details).
Experiment Setup
Experiments are conducted on an AMD Threadripper 3970x 32 core CPU with 256 GB RAM, and two NVIDIA A6000 GPU with 48 GB VRAM. We use Python (PyTorch deep learning framework) and Linux operating system.
Multimodal Baselines
We compare our proposed UFNet with four popular multimodal fusion approaches.
Majority Voting. The predictions from three different task-specific models are combined to generate a single prediction. The final prediction is the class (PD or Non-PD) agreed upon by the majority (i.e., two or more).
Neural Late Fusion. The logit scores from the task-specific models are input to a shallow neural network trained to predict a binary class, similar to logistic regression or other ensembling methods.
Early Fusion Baseline. Features from all three tasks are concatenated and input to a shallow neural network.
Hybrid Fusion Baseline. Task-specific features and prediction scores (logits) are both provided as input to a shallow neural network. This network leverages both the input features and prediction scores, combining the strengths of early and late fusion approaches.
Ablation Models. UFNet uses task-specific features and prediction scores, making it a hybrid fusion approach. Additionally, we analyze the effect of removing task-specific predictions from UFNet, implementing an early fusion approach. We also examine the impact of withholding predictions for both early and hybrid versions of the UFNet model.
Model Selection and Performance Reporting
The model with the highest AUROC on the validation set is selected for the reported experiments. After selection, each model is run with 30 different random seeds, and evaluated on the test set. The average and 95% confidence interval of each performance metric are reported. Due to dataset imbalance, we report accuracy, balanced accuracy, $F_{1}$ score, AUROC, and AUPRC to compare the models’ performance. Specificity, sensitivity, positive predictive value (PPV), and negative predictive value (NPV) of UFNet and other baselines are also reported. Coverage (of cases with a prediction) is provided when uncertain predictions are withheld. Significant differences mean no overlap in 95% confidence intervals; non-inferiority means overlap.
For patient safety, it is critical to ensure that the prediction scores approximate the true probability of the positive class. Therefore, we evaluate model calibration using expected calibration error (ECE) (Nixon et al. 2019) and Brier score (Rufibach 2010). Misclassification rates across male and female subgroups are compared using a two-sample $Z$-test for proportions to analyze potential bias. Model performance across white and non-white participants is compared using Fisher’s exact test due to the small number of non-white participants available in the test set. We also report average misclassification rates and confidence intervals across different age groups.
Results
Task-specific Model Performance
Task Accuracy
Balanced
Accuracy
$F_{1}$ score AUROC AUPRC
Finger-tapping
w/o MC-dropout $72.0\pm 0.9$ $69.0\pm 0.9$ $60.2\pm 1.3$ $73.9\pm 0.9$ $58.9\pm 1.0$
w/ MC-dropout $73.1\pm 0.7$ $70.1\pm 0.7$ $61.7\pm 0.9$ $74.9\pm 0.7$ $58.1\pm 0.9$
Smile
w/o MC-dropout $75.6\pm 0.2$ $72.2\pm 0.2$ $64.3\pm 0.3$ $83.2\pm 0.1$ $64.9\pm 0.2$
w/ MC-dropout $77.6\pm 0.2$ $74.4\pm 0.2$ $67.5\pm 0.3$ $83.6\pm 0.1$ $65.4\pm 0.1$
Speech
w/o MC-dropout $84.5\pm 0.3$ $82.5\pm 0.4$ $71.7\pm 0.4$ $\mathbf{89.4\pm 0.2}$ $\mathbf{81.9\pm 0.4}$
w/ MC-dropout $\mathbf{85.1\pm 0.2}$ $\mathbf{83.8\pm 0.5}$ $\mathbf{72.1\pm 0.6}$ $87.8\pm 0.1$ $80.7\pm 0.3$
Table 2: Performance of task-specific models with (w/) and without (w/o) MC-dropout (in percentage). Underlined metrics denote significantly better performance (within the same task), while bold metrics show overall best performance.
Among the three standardized tasks, the pangram utterance (speech) is the most accurate for classifying individuals with and without PD. Using only speech features, the shallow neural network achieved $84.5\pm 0.3\%$ accuracy and $89.4\pm 0.2\%$ AUROC. Finger-tapping had the least accuracy for PD detection. Models trained on right-hand tapping features performed better than those trained on left-hand features. However, the model trained on concatenated features from both hands significantly improved the $F_{1}$ score while being non-inferior in other metrics (see Technical Appendix B for details). For the remaining analysis, the finger-tapping model refers to the one trained on both-hand features. The model trained on smile features performed better than the finger-tapping task but worse than the speech task.
Effect of MC-dropout
MC-dropout significantly boosted the performance of the smile model in all metrics. For the speech model, while MC dropout improved accuracy and balanced accuracy, it decreased AUROC and AUPRC. The performance of the finger-tapping model was not notably affected. Due to the additional benefits of MC dropout in modeling prediction uncertainty, the multi-task models (UFNet and other baselines) were trained with MC-dropout unless specified otherwise. Detailed performance metrics of the task-specific neural networks along with the effect of MC dropout are reported in Table 2.
Effect of multi-task combinations
Combining multiple tasks using the proposed UFNet model enhanced performance (Table 3). For instance, the AUPRC scores of the multi-task models were significantly better than the corresponding single-task scores. Although the finger-tapping task alone was the weakest for detecting PD, its features may have complemented other task features, resulting in significant improvements in most metrics. Notably, combining all three tasks significantly improved all reported metrics, achieving an accuracy of $87.3\pm 0.3\%$, AUROC of $92.8\pm 0.2\%$, and an $F_{1}$ score of $81.0\pm 0.6\%$.
Task Combination Accuracy
Balanced
Accuracy
$\mathbf{F_{1}}$ score AUROC AUPRC
All three tasks $\mathbf{87.3\pm 0.4}$ $\mathbf{86.4\pm 0.4}$ $\mathbf{81.0\pm 0.6}$ $\mathbf{92.8\pm 0.2}$ $\mathbf{86.3\pm 0.5}$
Finger-tapping + Smile $78.0\pm 0.8$ $75.1\pm 1.0$ $65.6\pm 1.7$ $84.8\pm 0.5$ $74.5\pm 0.7$
Finger-tapping + Speech $84.1\pm 0.3$ $82.4\pm 0.3$ $77.3\pm 0.4$ $91.4\pm 0.2$ $84.5\pm 0.3$
Smile + Speech $85.2\pm 0.3$ $82.8\pm 0.5$ $75.0\pm 0.4$ $91.2\pm 0.1$ $82.2\pm 0.5$
Table 3: Performance of models (as percentages) trained on different combinations of the standardized tasks. The bold metrics denote best performance.
Model Accuracy
Balanced
Accuracy
AUROC AUPRC $\mathbf{F_{1}}$ score
PPV
NPV
Sensitivity
Specificity Coverage
Majority Voting 85.3 83.9 89.6 78.0 78.2 80.0 87.9 76.5 89.9 -
Neural Late Fusion $84.1\pm 0.4$ $81.3\pm 4.8$ $91.7\pm 2.2$ $\mathbf{86.7\pm 3.1}$ $73.2\pm 8.3$ $73.5\pm 7.5$ $89.2\pm 3.3$ $76.3\pm 9.4$ $88.2\pm 2.7$
Early Fusion Baseline $83.6\pm 0.6$ $81.8\pm 0.7$ $91.0\pm 0.2$ $85.8\pm 0.4$ $76.7\pm 0.7$ $75.4\pm 1.1$ $88.3\pm 0.4$ $78.1\pm 0.9$ $86.5\pm 0.8$
Hybrid Fusion Baseline $84.1\pm 0.3$ $82.4\pm 0.4$ $91.4\pm 0.2$ $86.5\pm 0.3$ $77.3\pm 0.4$ $76.2\pm 0.7$ $88.5\pm 0.3$ $78.6\pm 0.6$ $87.0\pm 0.6$
UFNet - Early Fusion $86.7\pm 0.5$ $85.8\pm 0.5$ $92.7\pm 0.3$ $86.2\pm 0.5$ $79.9\pm 0.8$ $83.3\pm 0.7$ $88.3\pm 0.6$ $76.9\pm 1.4$ $91.9\pm 0.4$
UFNet - Early Fusion
(withhold uncertain preds)
$87.5\pm 0.4$ $86.5\pm 0.5$ $92.9\pm 0.3$ $86.4\pm 0.7$ $80.7\pm 0.8$ $84.0\pm 0.7$ $89.1\pm 0.6$ $77.7\pm 1.4$ $92.4\pm 0.4$ $97.4\pm 0.3$
UFNet - Hybrid Fusion $87.3\pm 0.4$ $86.4\pm 0.4$ $92.8\pm 0.2$ $86.3\pm 0.5$ $81.0\pm 0.6$ $83.8\pm 0.5$ $89.0\pm 0.4$ $78.4\pm 1.0$ $92.0\pm 0.3$
UFNet - Hybrid Fusion
(withhold uncertain preds)
$\mathbf{88.0\pm 0.3}$ $\mathbf{87.1\pm 0.3}$ $\mathbf{93.0\pm 0.2}$ $86.5\pm 0.5$ $\mathbf{81.8\pm 0.6}$ $\mathbf{84.6\pm 0.5}$ $\mathbf{89.7\pm 0.4}$ $\mathbf{79.3\pm 0.9}$ $\mathbf{92.6\pm 0.2}$ $97.8\pm 0.3$
Table 4: Comparison of UFNet performance against multimodal baselines. The underlined metrics indicate significantly better performance compared to all four baselines. Bold metrics indicate overall best performance across all choices. All scores should be interpreted as percentages (%). Confidence interval is not reported for majority voting since it does not involve any randomness.
(a) ROC curve.
(b) Model calibration.
Figure 2: ROC and calibration curves showing the performance of the best models (with different random seeds). The shaded regions and the error bars represent 95% confidence intervals.
Comparison against baselines
The proposed uncertainty-calibrated fusion network (UFNet) outperformed all four baseline methods in most metrics (see Table 4). While the neural late fusion baseline achieved the best AUPRC, its average $F_{1}$ score is significantly lower compared to the other baselines and UFNet models. This baseline also demonstrates the least stability, with a large confidence interval for most metrics. UFNet significantly improves accuracy, balanced accuracy, PPV, and specificity over the baselines, with other metrics being non-inferior.
Ablation results
Including task-specific predictions as additional input (hybrid fusion) slightly enhanced model performance. Withholding uncertain predictions also boosted performance. With dropped uncertain predictions, the best UFNet (hybrid fusion) model achieved $88.0\pm 0.3\%$ accuracy, $87.1\pm 0.3\%$ balanced accuracy, $93.0\pm 0.2\%$ AUROC, and $81.8\pm 0.5\%$ $F_{1}$ score, outperforming all other models. It also achieves the best PPV ($84.6\pm 0.5\%$), NPV ($89.7\pm 0.4\%$), sensitivity ($79.3\pm 0.9\%$), and specificity ($92.6\pm 0.3\%$). However, the model can now predict for $97.8\pm 0.3\%$ of the data where it is certain enough. In addition, the expected calibration error (ECE) of the model is $5.4\pm 0.5\%$, and the Brier score is $0.097\pm 0.002$, indicating that the model predicted probability is aligned with true disease probability (Figure 2).
Figure 3: Misclassification rate of the best UFNet model across demographic subgroups. The analysis is done on the test set participants. Error bars demonstrate 95% confidence intervals.
Performance across demographic subgroups
No significant bias in model performance was observed based on sex or race in the test set (162 individuals). The average error (i.e., misclassification) rate across female participants ($n=85$) was $14.1\pm 7.4\%$, compared to $6.5\pm 5.5\%$ for male participants ($n=77$). This difference, though notable, is not statistically significant (p-value = 0.11). The error rate was $7.63\pm 4.79\%$ for white participants ($n=118$) and $5.56\pm 11.39\%$ for non-white participants ($n=18$). Based on Fisher’s exact test, this difference was also non-significant (Fisher’s odd ratio = 0.71, p-value = 1.0). However, the error rate varied notably based on age subgroups. The model performed well for individuals aged 50 to 80, while the error rate was higher for those aged 30 to 50 or over 80, likely due to under-representation of these age groups, as the dataset primarily consists of 50-80 year old individuals (77.1% of the entire dataset).
Discussions and Limitations
Video analysis provides an accessible, cost-effective, and convenient means of screening for PD, particularly benefiting individuals in remote areas or low-income countries with limited access to neurological care. Building on prior work demonstrating the feasibility of classifying PD symptoms from single video tasks (Rahman et al. 2021; Islam et al. 2023; Adnan et al. 2023), this study explores a combination of three tasks for a more holistic and generalizable approach, reflecting the multifaceted nature of PD symptoms. We developed models to classify PD cases with high accuracy, sensitivity, and specificity using webcam recordings of individuals performing finger tapping, speech, and smile tasks.
We carefully selected three proposed tasks for remote completion, considering feasibility and safety. These tasks assess bradykinesia, hypomimia, and speech impairment and can be safely performed at home without assistance. While gait analysis is common for evaluating PD (Li and Li 2022; Liu et al. 2022a), it presents logistical challenges and potential risks of falling for PD patients. In contrast, the finger-tapping task offers a safer alternative for motor assessment. Facial expressions and speech are assessed through natural conversations, but prompting natural conversations introduces subjectivity and confounding factors, hindering machine learning model training with limited data. Mimicking a smile provides a standardized alternative, and the pangram utterance task reduces confounding factors in speech analysis. While sustained phonation (holding a vowel sound for as long as possible) is another speech assessment option (Tsanas et al. 2009; Vaiciukynas et al. 2017), its analysis becomes complicated due to inconsistencies across various recording devices. All selected tasks demonstrated reasonable predictive performance in differentiating individuals with and without PD, validating our task selection.
Models trained on all three tasks performed better than single-task models. Among single-task models, PD classification from the speech task performed the best. However, the UFNet model trained on all three tasks was our best-performing model, achieving the highest AUROC, accuracy, recall, and specificity. Withholding uncertain predictions further boosted the model’s performance, ensuring additional safeguards against potential miss-predictions. The proposed model also outperformed traditional data-efficient neural models for combining multiple modalities.
However, our work has limitations. The model performs consistently across sex and race subgroups, but accuracy drops for younger ($30-49$) and older (above 80) age groups. We find that participants from these age groups are under-represented in our dataset (see Technical Appendix A containing demographic summary), with over 75% of the participants falling within the 50-80 age range. This age bias likely explains the model’s more robust performance in this middle range. It’s also important to note that younger individuals are less likely to have PD, as it is more commonly associated with aging. Until a more balanced dataset is available, we recommend applying the tool only for individuals between 50 and 80 years old. Future work should prioritize recruiting younger and older participants.
Additionally, the thresholds used in the study could be customized based on individual preferences, as the decision threshold directly affects the model’s sensitivity and specificity. We used a common 0.5 threshold, but individual preferences for risk-benefit trade-offs might necessitate adjustments. For instance, some users might seek clinical evaluation even at lower probabilities (i.e., prefer high sensitivity), while others might wait for a higher likelihood before incurring healthcare costs (i.e., prefer high specificity). We chose to withhold predictions when the model is uncertain about the PD/non-PD classification. If the 95% confidence interval of the prediction contains both positive and negative classes, the model refrains from making a prediction. In the absence of MC-dropout, an alternative is interpreting the model prediction scores as its confidence (e.g., a model is more confident when the prediction score is $0.95$ than a score of $0.55$) and setting a suitable confidence level based on accuracy-coverage trade-off and participant preference. Other alternatives for withholding predictions can also be explored in the future.
Furthermore, as videos are primarily gathered in an unsupervised fashion, issues such as noncompliance with task instructions and various forms of noise are common. For example, during the finger-tapping task, many participants performed fewer than the required ten taps, often with their task-performing hand obscured from view. Background noise may distort speech features, and the presence of multiple individuals in the frame could compromise smile feature extraction. Integrating post-hoc quality assessment algorithms into the data collection framework could enhance data quality by identifying quality issues and prompting the user to re-record tasks if needed.
A comprehensive, large-scale dataset for Parkinson’s disease is currently unavailable due to the lack of substantial resources and clinical collaboration. This paper addresses this gap by introducing the largest multi-task video dataset in the literature and publicly releasing the anonymized features. Unfortunately, we could not gather multiple datasets of videos or extract similar features to compare model performance across different datasets since patient videos, being protected health information, are not publicly accessible. Making our dataset available can help bridge this gap, facilitating broader comparisons and furthering research in this area.
Conclusion
This study demonstrates the promising efficacy of machine learning models in distinguishing individuals with PD from those without PD, requiring only a computer equipped with a webcam, microphone, and internet connection. Given the shared characteristics and nuanced distinctions among movement disorders such as Parkinson’s disease, Huntington’s disease, ataxia, and Progressive Supranuclear Palsy, these findings hold significant implications. We hope the promising initial results from this research will pave the way for extending tele-neurology applications to encompass a broader range of movement disorders.